40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 16 19:48:52 2016, Lydia, Update, General, Alignment status 
    Reply  Sat Sep 17 02:00:23 2016, rana, Update, SUS, Alignment status 
       Reply  Sat Sep 17 16:51:01 2016, Lydia, Update, SUS, Alignment status Screenshot_from_2016-09-17_16-45-00.png
          Reply  Sun Sep 18 16:18:05 2016, rana, Update, SUS, Alignment status 
             Reply  Mon Sep 19 11:11:43 2016, ericq, Update, SUS, Alignment status 
                Reply  Mon Sep 19 22:03:10 2016, ericq, Update, General, Further recovery progress 
                   Reply  Tue Sep 20 10:45:06 2016, rana, Update, General, Further recovery progress 
                      Reply  Wed Sep 21 01:08:02 2016, ericq, Update, General, Further recovery progress 
                         Reply  Wed Sep 21 09:04:57 2016, Steve, Update, General, 8 hours recovery progress 8hrs.png
                            Reply  Thu Sep 22 15:44:21 2016, Steve, Update, General, ITMX magnets are stucked again StuckAgain.png
                               Reply  Thu Sep 22 20:01:47 2016, Lydia, Update, General, ITMX freed, all optics kicked 
                                  Reply  Thu Sep 22 22:52:08 2016, ericq, Update, General, Damping found to be on 
Message ID: 12500     Entry time: Fri Sep 16 19:48:52 2016     Reply to this: 12501
Author: Lydia 
Type: Update 
Category: General 
Subject: Alignment status 

Today the Y arm was locking fine. The alignment had drifted somewhat so I ran the dither and TRY returned to ~0.8. However, the mode cleaner has been somewhat unstable. It locked many times but usually for only a few minutes. Maybe the alignment or autolocker needs to be adjusted, but I didn't change anything other than playing with the gain sliders (which didn't seem to make it either better or worse).

ITMX is still stuck.

ELOG V3.1.3-